OpenStudy (anonymous):

A box contains 3 bags of corn chips, 3 bags of potato chips, 3 bags of peanuts, and 3 bags of pretzels. One bag is chosen at random. What is the probability that it is a bag of peanuts? A. 1/2 B. 1/12 C. 1/4 D. 1/3

OpenStudy (lucaz):

you have a total of 12 bags, and 3 of them are bags of peanuts P(peanuts)=3/12=1/4

OpenStudy (lucaz):

ideia is (what you want) / (all you have)